If 4 times a number is increased by 1/2 itself, the result is 144. What is the number
Answer by Boreal(15235)

Let x = the number.
4x is four times the number
1/2 of itself is 1/2 x.
4 times increased by half itself is 4x + (1/2)x
This equals 144.
4 (1/2) x=144; (9/2)x =144 ; x=144 * (2/9)=32
Check
32 *4=128
Add half of 32 (16) and get 144.